# Finite-Agent Stochastic Differential Games on Large Graphs: I. The Linear-Quadratic Case

### v1

@@ 1 Introduction
For a linear-quadratic game, numerically solving for the NE can always be transformed into numerically solving the associated coupled Riccati equations. This task is relatively straightforward due to the existence of numerous efficient numerical schemes for solving ODEs. However, solving for the NE of a general stochastic differential game is challenging, especially when the game is fully nonlinear (with controlled diffusion terms), involves a relatively large number of players, and/or has high-dimensional states. This complexity gives rise to fictitious play (FP) as an efficient method for numerically solving for the NE of an non-cooperative game in an iterative way, initially introduced by Brown [Bro49, Bro51]. The main idea behind FP is to recast the problem into a sequence of stochastic control problems faced by each player to be solved repeatedly. However, the convergence of FP to the NE is not guaranteed, even for tabular games [Sha63], which motivates research on the convergence of FP. In the context of continuous-time differential games, affirmative results include [CH17, BC18] for mean-field games and [Hu21, HHL22] for finite-player games.
Moreover, we emphasize that this paper focuses on directly tackling finite-agent games rather than using mean-field game approximations. Mean-field games (MFGs), first proposed by Lasry and Lions [LL07], provide asymptotic approximations to finite-agent games with mean-field interactions. MFGs have been extensively used to model human behaviors in various fields, including Aiyagari’s growth model [CD18] in economics, the systemic risk model [CFS15] in finance, and the cryptocurrency mining model [LRS24] in blockchain technology, among others. Our choice of directly addressing the finite population problem is due to several reasons. Despite the usefulness of mean-field games, their approximations to the finite population counterpart rely on the assumption of a large number of identical players, which does not always hold. The recently developed graphon MFGs [CH19, CH21, GCH23] address this issue by modeling the heterogeneous interactions through graphons, which are the limits of dense graphs. However, sparse social networks are also of significant research interest. In contrast, finite-player games can model heterogeneous interactions, sparse or dense, directly and can accommodate a relatively smaller number of players, which MFGs cannot. For further research on learning graphon mean-field games, we refer readers to [CK21, FCK23].
Main contribution. Our main contributions in this paper are as follows: (i) We propose a new finite-agent linear-quadratic game on graphs, aiming to provide a more realistic modeling of players with heterogeneous and interpretable interactions. This substantially extends the scope of existing research, such as [CFS15], which primarily focuses games on equally-weighted complete graphs, and [LS22] by incorporating information flow from neighboring players. (ii) With arbitrary graphs, we establish the convergence of fictitious play for solving the Markovian NE of our model under a smallness condition, regardless of the number of players. This improves the previous analysis [Hu21], which addresses the open-loop Nash equilibrium of games featuring homogeneous players on complete graphs, and also provides additional theoretical backing for the feasibility of FP. (iii) For vertex-transitive graphs, we construct a semi-explicit characterization of the NE. The semi-explicit characterization under our model is intricately linked to solving a matrix-valued ODE, which degenerates to a scalar ODE in a special case [LS22]. We investigate the existence and uniqueness of the solution to the matrix-valued ODE, which requires substantial effort to establish. (iv) The numerical experiments not only validate our theoretical results but also offer crucial insights into the relationship between various aspects of the game and the underlying graph structure.
@@ 2.1 The game setup
[…]
@@
$f^{i}(t,x,\alpha)=\frac{1}{2}(\alpha)^{2}-q\alpha\left[\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}x^{j}-x^{i}\right]+\frac{\varepsilon}{2}\left[\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}x^{j}-x^{i}\right]^{2},$ (2.5)
$g^{i}(x)=\frac{c}{2}\left[\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}x^{j}-x^{i}\right]^{2}.$ (2.6)
The parameters are in the range $q>0$, $c>0$, $\varepsilon>0$, and the condition $q^{2}\leq\varepsilon$ is required to ensure the joint convexity of $f^{i}$ and the well-posedness of the problem.
@@
$\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}X^{j}_{t}=\frac{1}{d_{v_{i}}}\sum_{j:v_{j}\sim v_{i}}\sqrt{\frac{d_{v_{i}}}{d_{v_{j}}}}X^{j}_{t},$ (2.7)
which is a weighted state average among all the players in the neighborhood of $v_{i}$. We interpret the square root weight in equation (2.7) from two perspectives.
The first perspective comes from the connection with graph Laplacian $L$. Let $\{e_{1},\ldots,e_{N}\}$ be the standard basis of $\mathbb{R}^{N}$. From Definition 2.1, it is clear that
$\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}X^{j}_{t}-X^{i}_{t}=-e_{i}^{T}LX_{t}.$ (2.8)
This concise representation facilitates the connection between the graph structure and various aspects of stochastic differential games on graphs in later discussions.
The second perspective comes from considering the random walk on graphs as a process for transmitting substances or diverging information across a network. A standard result for the random walk on graph, as outlined in [Chu97], is the stationary distribution $\pi_{i}=d_{v_{i}}/\text{vol}(G)$, where the volume of the graph $G$ is defined as $\text{vol}(G)=\sum_{i=1}^{N}d_{v_{i}}$, serving as a normalization factor. Let $p_{ij}$ denote the one-step transition probability from $v_{i}$ to $v_{j}$. The detailed balance equation of this random walk on graphs is given by $d_{v_{i}}/d_{v_{j}}=p_{ji}/p_{ij}$. Now, consider the symmetry constraint that the transmission/divergence of substance/information along each undirected edge is symmetric in both directions, i.e., if $p_{ji}=h(j,i)$ for some function $h$, then $p_{ij}=h(i,j)$ must also hold. Under the detailed balance equation and the symmetry constraint along undirected edges, $p_{ji}=\sqrt{d_{v_{i}}/d_{v_{j}}}$ provides a natural way to construct the weights for the mean reversion level of the model.
Note that if $G$ is a regular graph, meaning all vertices have the same degree, the mean reversion level becomes the arithmetic average of the states of all players in the neighborhood of $v_{i}$.
@@
In a stochastic differential game, the information set on which each player makes its decision must be specified. The difference in the information set results in different notions of strategy, e.g., open-loop strategy, closed-loop strategy, Markovian strategy, etc. In this paper, our discussion is restricted to the Markovian strategy, i.e., $\alpha^{i}_{t}=\phi^{i}(t,X_{t})$ for some deterministic feedback function $\phi^{i}:[0,T]\times\mathbb{R}^{N}\to\mathbb{R}$ of player $i$. To match the restrictions in the definition of $\mathscr{A}$ (cf. (2.3)), the feedback function $\phi^{i}$ shall be Borel measurable such that $\sup_{(t,x)\in[0,T]\times\mathbb{R}^{N}}\frac{|\phi^{i}(t,x)|}{1+\|x\|}<\infty$. In the setting of a competitive game, the Nash equilibrium defined below is of general interest.
@@
In simple words, an Nash equilibrium is a collection of strategies from which any player does not have the motivation to deviate, given that all other players stick to their Nash equilibrium strategies.
@@ 2.2 The associated HJB system
$\partial_{t}v^{i}+\inf_{\alpha^{i}}\left\{\sum_{k=1}^{N}(\alpha^{k}-ae_{k}^{\operatorname{T}}Lx)\partial_{x^{k}}v^{i}+\frac{\sigma^{2}}{2}\sum_{k=1}^{N}\partial_{x^{k}x^{k}}v^{i}+\frac{1}{2}(\alpha^{i})^{2}+q\alpha^{i}e_{i}^{\operatorname{T}}Lx+\frac{\varepsilon}{2}(e_{i}^{\operatorname{T}}Lx)^{2}\right\}=0,$ (2.10)
@@
where $F^{i}:[0,T]\to\mathbb{S}^{N\times N}$ and $h^{i}:[0,T]\to\mathbb{R}$ are both deterministic functions in time. Plugging equations (2.11)–(2.12) back into equation (2.10), and collecting the coefficients of $x$ yield a Riccati equation for $F^{i}$:
$\dot{F}^{i}_{t}+\sum_{k=1}^{N}[-(a+q)L-F^{k}_{t}]e_{k}e_{k}^{\operatorname{T}}F^{i}_{t}+\sum_{k=1}^{N}F^{i}_{t}e_{k}e_{k}^{\operatorname{T}}[-(a+q)L-F^{k}_{t}]\\ +(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L+F^{i}_{t}e_{i}e_{i}^{\operatorname{T}}F^{i}_{t}=0,\quad F^{i}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L.$ (2.13)
@@
This equilibrium depends on $F^{i}$ but not on $h^{i}$. Consequently, the rest of the paper will focus on analyzing the Riccati system (2.13).
@@ 3 The Equilibrium via Fictitious Play
The convergence of FP to the NE, however, is not guaranteed, even for tabular games [Sha63], which motivates research on the convergence of FP. In the context of continuous-time differential games, affirmative answers include [CH17, BC18] for mean-field games and [Hu21, HHL22] for finite-player games, but none of them take into account the graph structure. In this section, we provide proof for the convergence of FP for the generic model described in Section 2.1. This result contributes in two significant ways. Firstly, it validates the effectiveness of FP for solving stochastic differential games on graphs. Secondly, it elucidates the connection between the graph structure and the convergence of FP.
@@ 3.1 The mathematical formulation
To prepare for stating the convergence result in later contexts, we describe the procedure of FP in mathematical terms. Firstly, rewrite the state dynamics and the cost functionals in terms of the graph Laplacian $L$. Using equation (2.8), the state dynamics (2.2) is rewritten as
@@
For any object $A$, $A^{i,k}$ refers to the object associated with player $i\in[N]$ at FP stage $k\in\mathbb{N}$. Denote by $A^{k}$ the collection of objects across all the players at FP stage $k$. In the following context, $i$ and $j$ stand for the player indices, while $k$ represents the FP stage index.
Without loss of generality, let us consider the procedure of FP for player $i$. Assume all players have their strategies at stage $k$ determined through a feedback functions $\phi^{i,k}$ that $\alpha^{i,k}_{t}=\phi^{i,k}(t,X_{t}^{k})$, $i\in[N]$. At stage $k+1$, player $i$ optimizes her strategy to get a new feedback function $\phi^{i,k+1}$ based on the environment at stage $k$ while assuming all the other players stick to their feedback functions at stage $k$, i.e., player $j\neq i$ is using feedback function $\phi^{j,k}$. As a result, the stochastic control problem faced by player $i$ has state dynamics as
@@
$\partial_{t}v^{i,k+1}+\inf_{\alpha}\Big{\{}(\alpha-ae_{i}^{\operatorname{T}}Lx)\partial_{x^{i}}v^{i,k+1}+\sum_{j\neq i}(\phi^{j,k}(t,x)-ae_{j}^{\operatorname{T}}Lx)\partial_{x^{j}}v^{i,k+1}+\frac{\sigma^{2}}{2}\sum_{j=1}^{N}\partial_{x^{j}x^{j}}v^{i,k+1}\\ +\frac{1}{2}\alpha^{2}+q\alpha e_{i}^{\operatorname{T}}Lx+\frac{\varepsilon}{2}(e_{i}^{\operatorname{T}}Lx)^{2}\Big{\}}=0,\quad v^{i,k+1}(T,x)=\frac{c}{2}(e_{i}^{\operatorname{T}}Lx)^{2}.$ (3.4)
The first-order condition in the above equation gives
@@
The term $\phi^{j,k}(t,x)$ in the above equation hinders us from reducing the HJB equation to a Riccati equation directly. To proceed, we make the following assumption on the initial strategies $\phi^{i,0}$, for all players $i\in[N]$.
@@
Assume that for every player $i$, $i\in[N]$, FP starts with $\phi^{i,0}$ that is of the form $\phi^{i,0}(t,x)=(\varphi^{i,0}_{t})^{\operatorname{T}}x$, for some $\varphi^{i,0}:[0,T]\to\mathbb{R}^{N}$.
@@
Under Assumption 3.2, the optimal strategy $\phi^{i,k}$ for player $i$ at FP stage $k$, can always be represented as $\phi^{i,k}(t,x)=(\varphi^{i,k}_{t})^{\operatorname{T}}x$ for some $\varphi^{i,k}:[0,T]\to\mathbb{R}^{N}$, for $i\in[N]$ and $k\in\mathbb{N}$.
@@
$\dot{F}^{i,k+1}_{t}-F^{i,k+1}_{t}e_{i}e_{i}^{\operatorname{T}}F^{i,k+1}_{t}+(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L-aLF^{i,k+1}_{t}-aF^{i,k+1}_{t}L+\sum_{j\neq i}F^{i,k+1}_{t}e_{j}(\varphi^{j,k}_{t})^{\operatorname{T}}\\ +\sum_{j\neq i}\varphi^{j,k}_{t}e_{j}^{\operatorname{T}}F^{i,k+1}_{t}-qLe_{i}e_{i}^{\operatorname{T}}F^{i,k+1}_{t}-qF^{i,k+1}_{t}e_{i}e_{i}^{\operatorname{T}}L=0,\quad F^{i,k+1}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L.$ (3.7)
Plugging the ansatz into equation (3.5) yields $\phi^{i,k+1}(t,x)=-e_{i}^{\operatorname{T}}(F^{i,k+1}_{t}+qL)x$. Therefore, $\forall i\in[N]$ and $\forall k\in\mathbb{N}$,
@@
Note that $\{F^{i,k+1}\}_{i\in[N]}$ and $\{F^{i,k}\}_{i\in[N]}$ are coupled through the term $\varphi^{j,k}_{t}$, represented as $\varphi^{j,k}_{t}=-(F^{j,k}_{t}+qL)e_{j}$ (cf. (3.8)). This yields the following recursive Riccati equation, which describes the update from $\{F^{i,k}\}_{i\in[N]}$ to $\{F^{i,k+1}\}_{i\in[N]}$:

### v2

@@ 1 Introduction
For a linear-quadratic game, numerically solving for the NE can always be transformed into numerically solving the associated coupled Riccati equations. This task is relatively straightforward due to the existence of numerous efficient numerical schemes for solving ODEs. However, solving for the NE of a general stochastic differential game is challenging, especially when the game is fully nonlinear (with controlled diffusion terms), involves a relatively large number of players, and/or has high-dimensional states. This complexity gives rise to fictitious play (FP) as an efficient method for numerically solving for the NE of a non-cooperative game in an iterative way, initially introduced by Brown [Bro49, Bro51]. The main idea behind FP is to recast the problem into a sequence of stochastic control problems faced by each player to be solved repeatedly. However, the convergence of FP to the NE is not guaranteed, even for tabular games [Sha63], which motivates research on the convergence of FP. In the context of continuous-time differential games, affirmative results include [CH17, BC18] for mean-field games and [Hu21, HHL22] for finite-player games.
A closely related research direction is graphon mean-field games (graphon MFGs), which extend classical MFGs by incorporating heterogeneous interactions through graphons—limiting objects representing the structure of dense networks [BWZ23, CH19, CH21, GCH23]. Classical MFGs [LL07] approximate finite-agent games in the limit of an infinite number of homogeneous players and have been widely used to model human behavior in various fields, including Aiyagari’s growth model in economics [CD18], systemic risk models in finance [CFS15], and cryptocurrency mining models in blockchain technology [LRS24]. While classical MFGs offer potential analytical traceability and dimension reduction in large-population settings, their validity depends on the assumption that all players are identical in distribution, which may not hold in many practical scenarios. Graphon MFGs address this limitation by modeling structured interactions, but they rely on the assumption that the underlying network is dense. However, many real-world networks, including social networks, are often sparse, where graphon-based approximations may be less effective. In contrast, our setting directly models finite-agent games, accommodating both sparse and dense interactions without requiring a large game size. For further research on learning graphon mean-field games, we refer readers to [CK21, FCK23].
In this paper, we propose a new finite-agent linear-quadratic game on graphs, aiming to provide a more realistic modeling of players with heterogeneous and interpretable interactions. More precisely, we consider an $N$-player game, where the dynamics of player $i$ are given by
$\mathrm{d}X^{i}_{t}=\left[a\left(\bar{X}_{t}^{i}-X^{i}_{t}\right)+\alpha^{i}_{t}\right]\,\mathrm{d}t+\sigma\,\mathrm{d}W^{i}_{t},\quad\forall i\in[N],$ (1.1)
where $\alpha^{i}$ is player $i$’s strategy and $\bar{X}_{t}^{i}:=\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}X^{j}_{t}$ denotes the local average. Here $d_{v_{i}}$ represents the degree of node $i$ in the graph, characterizing the connectivity among $N$ players. The dynamics exhibit a mean-reverting nature, where the reversion level is determined by the state average of neighboring players. One key motivation for this model is peer influence in social networks, where $X_{t}^{i}$ represents player $i$’s opinion and the dynamics reflect opinion conformity.
This framework extends existing research, including [CFS15], which primarily focuses on games on equally-weighted complete graphs (where $d_{v_{i}}\equiv N-1$, $\forall i$), and [LS22], which assumes $a=0$ thus does not include mean reversion. By incorporating local interactions through the graph structure, our model captures heterogeneous influences and provides a more general setting for strategic decision-making. For further details of the model, including motivation and interpretation, see Section 2.1, particularly Remark 2.3.
Main Contribution. In Section 3, we analyze games on arbitrary graphs. Using dynamic programming and a properly chosen ansatz, we translate the game into a system of $N$ coupled matrix-valued Riccati equations, whose solvability has not yet been fully characterized in the literature. We establish local existence and uniqueness of the solution under a smallness condition, regardless of the number of players $N$. This is a key improvement over previous results such as [PC79], where the existence condition depends explicitly on $N$. Our approach leverages the convergence of fictitious play to solve for the Markovian NE, which extends the previous analysis [Hu21] that only addresses the case of open-loop NE for games with homogeneous players on complete graphs. In addition, our work provides theoretical backing for the feasibility of FP.
In Section 4, we focus on games on vertex-transitive graphs. The added symmetry among players allows us to refine the solution obtained from the Riccati system and construct a semi-explicit characterization of the NE. This characterization is intricately linked to solving a matrix-valued ODE, whose existence and uniqueness are far from trivial. Our key results include local well-posedness (Theorem 4.6), established via the Picard-Lindelöf theorem, and global existence and uniqueness (Theorem 4.7), obtained through a concavity argument that provides an a priori upper bound.
For comparison, in [CFS15], the Riccati equations are scalar-valued and decoupled, allowing for explicit solutions. In [LS22], where $a=0$, the ODE degenerates to a scalar equation, whose well-posedness follows from classical theory.
@@ 2.1 The game setup
$\mathscr{A}:=\left\{\alpha:\alpha\ \text{is progressively measurable w.r.t.}\ \{\mathscr{F}_{t}\},\ \mathbb{E}\int_{0}^{T}|\alpha_{t}|^{2}\mathrm{d}t<\infty\right\},$ (2.3)
@@
$f^{i}(t,x,\alpha)$ $=\frac{1}{2}(\alpha)^{2}-q\alpha\Big{[}\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}x^{j}-x^{i}\Big{]}$
$+\frac{\varepsilon}{2}\Big{[}\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}x^{j}-x^{i}\Big{]}^{2},$ (2.5)
$g^{i}(x)$ $=\frac{c}{2}\Big{[}\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}x^{j}-x^{i}\Big{]}^{2}.$ (2.6)
The parameters are in the range $q\geq 0$, $\varepsilon\geq 0$, $c>0$, and the condition $q^{2}\leq\varepsilon$ is required to ensure the non-negativity of $f^{i}$ and the joint convexity of $f^{i}$ in $(x,\alpha)$, which guarantees the well-posedness of the problem.
@@
$\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}X^{j}_{t}-X^{i}_{t}=-e_{i}^{T}LX_{t},$ (2.7)
where $\{e_{1},\ldots,e_{N}\}$ is the standard basis of $\mathbb{R}^{N}$. This concise representation facilitates the connection between the graph structure and various aspects of stochastic differential games on graphs in later discussions.
One motivation for studying the dynamics (2.2) comes from modeling peer influence in a given social network $G$. Here, $X^{i}_{t}$ represents the opinion of player $i$ on a certain topic, and the interactions occur only between neighboring nodes (friends). Under opinion conformity (where an individual’s opinion tends to align with those of others), the opinion dynamics naturally follow a mean-reversion process within each graph neighborhood. Players can actively adjust their opinions via taking actions $\alpha^{i}_{t}$, incurring costs for deviations and exertion of effort. The cost functionals (2.1)–(2.6) encode the trade-off: each player seeks to conform to the average opinion of their immediate neighbors while minimizing adjustment costs.
We note that different definitions of the graph Laplacian exist in the graph theory literature. The unnormalized graph Laplacian $L^{\text{UN}}$ is a symmetric matrix with real eigenvalues [GR01]. However, its eigenvalues admit possible dependence on $N$, and its spectral properties are dominated by high-degree graph vertices. Among normalized graph Laplacian, one commonly used variant is the random walk normalized graph Laplacian $L^{\text{RW}}$ defined in (2.8), which is motivated by the transition probability of random walks on graphs:
$L^{\text{RW}}_{ij}:=\begin{cases}1&\text{if}\ i=j\\ -\frac{1}{d_{v_{i}}}&\text{if}\ i\neq j\ \text{and}\ v_{i}\sim v_{j}\\ 0&\text{else}\end{cases},\ \forall i,j\in[N].$ (2.8)
The mean reversion level associated with $L^{\text{RW}}$ corresponds to the arithmetic average within the graph neighborhood:
$\frac{1}{d_{v_{i}}}\sum_{j:v_{j}\sim v_{i}}X^{j}_{t}-X^{i}_{t}=-e_{i}^{T}L^{\text{RW}}X_{t}.$ (2.9)
Nevertheless, the asymmetry of $L^{\text{RW}}$ on general graphs introduces technical challenges and complicates theoretical analysis. To overcome these difficulties, we base our discussion on the symmetric normalized graph Laplacian $L$, as defined in Definition 2.1. A straightforward calculation shows that $L^{\text{RW}}$ and $L$ are similar matrices, related by $L^{\text{RW}}=D^{-\frac{1}{2}}LD^{\frac{1}{2}}$, where $D:=\mathrm{diag}(d_{v_{1}},\ldots,d_{v_{N}})$ is the diagonal degree matrix. Thus, $L$ can be understood as a symmetric version of $L^{\text{RW}}$ that maintains all its spectral information.
While the above discussion motivates our Laplacian-based mean reversion formulation in equation (2.7), it is worth noting that the results in Section 3 extend to a broader class of averaging schemes beyond Laplacian-based ones, i.e., for different choices of symmetric matrices $L$, as long as $L$ has a spectrum bounded by a constant. However, in Section 4, where we focus on games over vertex-transitive graphs, a high degree of symmetry is required, making the normalized Laplacian-based formulation particularly natural. Therefore, we choose to adopt this specific formulation (2.7) throughout the following context.
Note that if $G$ is a regular graph, meaning all vertices have the same degree, the mean reversion level in (2.7) becomes the arithmetic average of the states of all players in the neighborhood of $v_{i}$.
@@
In a stochastic differential game, the information set based on which each player makes its decision must be specified. The difference in the information set results in different notions of strategy, e.g., open-loop strategy, closed-loop strategy, Markovian strategy, etc. In this paper, our discussion is restricted to the Markovian strategy, i.e., $\alpha^{i}_{t}=\phi^{i}(t,X_{t})$ for some deterministic feedback function $\phi^{i}:[0,T]\times\mathbb{R}^{N}\to\mathbb{R}$ of player $i$. To match the restrictions in the definition of $\mathscr{A}$ (cf. (2.3)), the feedback function $\phi^{i}$ shall be Borel measurable such that $\sup_{(t,x)\in[0,T]\times\mathbb{R}^{N}}\frac{|\phi^{i}(t,x)|}{1+\|x\|}<\infty$. In the setting of a competitive game, the Nash equilibrium defined below is of general interest.
@@
In simple words, a Nash equilibrium is a collection of strategies from which any player does not have the motivation to deviate, given that all other players stick to their Nash equilibrium strategies.
@@ 2.2 The associated HJB system
$\partial_{t}v^{i}+\inf_{\alpha^{i}}\Big{\{}\sum_{j=1}^{N}(\alpha^{j}-ae_{j}^{\operatorname{T}}Lx)\partial_{x^{j}}v^{i}+\frac{\sigma^{2}}{2}\sum_{j=1}^{N}\partial_{x^{j}x^{j}}v^{i}\\ +\frac{1}{2}(\alpha^{i})^{2}+q\alpha^{i}e_{i}^{\operatorname{T}}Lx+\frac{\varepsilon}{2}(e_{i}^{\operatorname{T}}Lx)^{2}\Big{\}}=0,$ (2.11)
@@
where $F^{i}:[0,T]\to\mathbb{S}^{N\times N}$ and $h^{i}:[0,T]\to\mathbb{R}$ are both deterministic and measurable functions in time. Plugging equations (2.12)–(2.13) back into equation (2.11), and collecting the coefficients of $x$ yield a Riccati equation for $F^{i}$:
$\dot{F}^{i}_{t}+\sum_{j=1}^{N}[-(a+q)L-F^{j}_{t}]e_{j}e_{j}^{\operatorname{T}}F^{i}_{t}+\sum_{j=1}^{N}F^{i}_{t}e_{j}e_{j}^{\operatorname{T}}[-(a+q)L-F^{j}_{t}]\\ +(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L+F^{i}_{t}e_{i}e_{i}^{\operatorname{T}}F^{i}_{t}=0,\quad F^{i}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L.$ (2.14)
@@
This equilibrium strategy depends on $F^{i}$ but not on $h^{i}$. Consequently, the rest of the paper will focus on analyzing the Riccati system (2.14).
@@ 3 The Equilibrium via Fictitious Play
The convergence of FP to the NE, however, is not guaranteed, even for tabular games [Sha63], which motivates research on the convergence of FP. In the context of continuous-time differential games, affirmative answers include [CH17, BC18] for mean-field games and [Hu21, HHL22] for finite-player games, but none of them take into account the graph structure. In this section, we provide proof for the convergence of FP for the generic model described in Section 2.1. This result contributes in three significant ways. Firstly, it provides a method for constructing the NE under conditions that are independent of the game size $N$. Secondly, it validates the effectiveness of FP for solving stochastic differential games on graphs. Thirdly, it elucidates the connection between the graph structure and the convergence of FP.
@@ 3.1 The mathematical formulation
To prepare for stating the convergence result in later contexts, we describe the procedure of FP in mathematical terms. Firstly, rewrite the state dynamics and the cost functionals in terms of the graph Laplacian $L$. Using equation (2.7), the state dynamics (2.2) can be rewritten as
@@
For any object $A$, $A^{i,k}$ refers to the object associated with player $i$ at FP stage $k\in\mathbb{N}$. Denote by $A^{k}$ the collection of objects across all the players at FP stage $k$. In the following context, $i$ and $j$ stand for the player indices, while $k$ represents the FP stage index.
Without loss of generality, let us consider the procedure of FP for player $i$. Assume all players have their strategies at stage $k$ determined through feedback functions $\phi^{j,k}$ that $\alpha^{j,k}_{t}=\phi^{j,k}(t,X_{t}^{k})$, $\forall j\in[N]$. At stage $k+1$, player $i$ optimizes her strategy to get a new feedback function $\phi^{i,k+1}$ based on the environment at stage $k$ while assuming all the other players stick to their feedback functions at stage $k$, i.e., player $j\neq i$ is using feedback function $\phi^{j,k}$. As a result, the stochastic control problem faced by player $i$ has state dynamics as
@@
Remark 3.2 (Variants of FP algorithms).
The original fictitious play (FP) [Bro49, Bro51] updates strategies based on the best response to the empirical distribution of opponents’ past actions. It has been shown to be an effective method for converging to a NE in normal-form games [Rob51, Hah99, Ber05].
In mean-field games (MFGs), FP has been adapted to update strategies based on the average density/measure of past population distributions [BC18, CH17, MH21]. An alternative approach, which we adopt here, updates strategies using only the most recent interaction. This method has been employed in both MFGs [HHL24] and finite-player games [Hu21, HH20, HHL22, XBH+22]. The primary motivation behind this choice is computational and due to memory efficiency—averaging strategies is often infeasible when strategies are parameterized by neural networks, since computing empirical averages introduces additional computational costs.
All the FP variants mentioned above follow simultaneous FP, meaning that the $(k+1)^{th}$ strategies for all players can be computed at the same time. In contrast, alternating FP updates strategies sequentially in a specific order. A thorough discussion of these variants can be found in [Hu21, Sections 5.2–5.4].
@@
$\partial_{t}v^{i,k+1}+\inf_{\alpha}\Big{\{}(\alpha-ae_{i}^{\operatorname{T}}Lx)\partial_{x^{i}}v^{i,k+1}+\sum_{j\neq i}(\phi^{j,k}(t,x)-ae_{j}^{\operatorname{T}}Lx)\partial_{x^{j}}v^{i,k+1}\\ +\frac{\sigma^{2}}{2}\sum_{j=1}^{N}\partial_{x^{j}x^{j}}v^{i,k+1}+\frac{1}{2}\alpha^{2}+q\alpha e_{i}^{\operatorname{T}}Lx+\frac{\varepsilon}{2}(e_{i}^{\operatorname{T}}Lx)^{2}\Big{\}}=0,$ (3.4)
with terminal condition $v^{i,k+1}(T,x)=\frac{c}{2}(e_{i}^{\operatorname{T}}Lx)^{2}$. The first-order condition in the above equation gives
@@
The term $\phi^{j,k}(t,x)$ in the above equation hinders us from reducing the HJB equation to a Riccati equation directly. To proceed, we make the following assumption on the initial strategies $\phi^{j,0}$, for all players $j\in[N]$.
@@
Assume that for every player $j$, $j\in[N]$, FP starts with $\phi^{j,0}$ that is of the form $\phi^{j,0}(t,x)=(\varphi^{j,0}_{t})^{\operatorname{T}}x$, for some $\varphi^{j,0}:[0,T]\to\mathbb{R}^{N}$.
@@
Under Assumption 3.3, the optimal strategy $\phi^{i,k}$ for player $i$ at FP stage $k$, can always be represented as $\phi^{i,k}(t,x)=(\varphi^{i,k}_{t})^{\operatorname{T}}x$ for some $\varphi^{i,k}:[0,T]\to\mathbb{R}^{N}$, for any $k\in\mathbb{N}$.
@@
$\dot{F}^{i,k+1}_{t}-F^{i,k+1}_{t}e_{i}e_{i}^{\operatorname{T}}F^{i,k+1}_{t}+(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L-aLF^{i,k+1}_{t}-aF^{i,k+1}_{t}L\\ +\sum_{j\neq i}F^{i,k+1}_{t}e_{j}(\varphi^{j,k}_{t})^{\operatorname{T}}+\sum_{j\neq i}\varphi^{j,k}_{t}e_{j}^{\operatorname{T}}F^{i,k+1}_{t}-qLe_{i}e_{i}^{\operatorname{T}}F^{i,k+1}_{t}-qF^{i,k+1}_{t}e_{i}e_{i}^{\operatorname{T}}L=0,$ (3.7)
with terminal condition $F^{i,k+1}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L$.
Plugging the ansatz into equation (3.5) yields $\phi^{i,k+1}(t,x)=-e_{i}^{\operatorname{T}}(F^{i,k+1}_{t}+qL)x$. Therefore, $\forall k\in\mathbb{N}$,
@@
Note that $\{F^{j,k+1}\}_{j\in[N]}$ and $\{F^{j,k}\}_{j\in[N]}$ are coupled through the term $\varphi^{j,k}_{t}$, represented as $\varphi^{j,k}_{t}=-(F^{j,k}_{t}+qL)e_{j}$ (cf. (3.8)). This yields the following recursive Riccati equations, which describe the update from $\{F^{j,k}\}_{j\in[N]}$ to $F^{i,k+1}$: